OpenStudy (anonymous):

What is the y-coordinate of the vertex of y = |–3x + 6| – 2 ? i got 0 but i'm not 100% sure

OpenStudy (campbell_st):

well not quite right... the y intercept occurs when x = 0 so substitute x = 0 and solve

OpenStudy (campbell_st):

and the vertex.... well what values makes -3x + 6 = 0 and the y value is found by letting |-3x + 6| = 0 and whats left

OpenStudy (anonymous):

so yea the y coordinate of the vertex, is -2

OpenStudy (anonymous):

memorize your transformations and basic functions :)

OpenStudy (campbell_st):

or just use a table of values.... x = -2, -1, 0, 1, 2, 3, 4, etc... makes it easy to see what the graph looks like